An UBER dispatcher finds that successive requests for UBER service are exponentially distributed with a mean time between requests of 10.50 minutes.

The dispatcher must disconnect the request receipt system for the next 3 minutes to reboot the system’s monitor.

What is the probability someone will attempt to contact this dispatcher while the system is disconnected for the monitor repair and reboot?
Answer #1

Ans:

mean=10.50 minutes

P(T<=t)=1-exp(-t/10.50)

P(T<=3)=1-exp(-3/10.50)=0.2485